Make New Entry,
Make Followup Entry
User name V. Sulkosky
Log entry time 19:32:39 on October 15, 2009
Entry number 295095
This entry is a followup to: 294939
keyword=Run Coordinator summary: October 14, Swing shift to 15, Day shift
RC report (Oct. 14 Swing shift ~ Oct. 15 Day shift)
Wednesday, Oct. 14, (Swing)
- 16:56 Beam back after injector bunchlength and harp scan
- 17:00 Resume production at 100 uA.
- Two online analysis crashes during swing shift at 18:40 and 20:25.
- Since beam is back, the lumi asymmetries are about 1500 ppm
probably due to fluctuations in coolant supply temperature.
- 20:25 Reduce current to 95 uA, lumi widths improve to 1200 ppm.
- 23:25 Begin 15 uA running for Compton electron detector and
Thursday, Oct. 15, (Owl and Day)
- 00:00 Production running at 100 uA.
- 03:00 Reduced current to 96 uA since lumi widths were around 1700
- 02:35 Compton rates flat-lined due to DAQ problem.
- 03:42 Beam IHWP inserted and began slug 24.
- 04;30 Trim rack bulk PS tripped off.
- 05:45 MCC informs shift leader that Moller dipole was left on
after the last Moller on Tuesday.
Moller dipole was turned off from shift leader's request.
- Detector linearity tests while beam is off.
- 08:00 MCC ready to send beam back to Hall A.
- 08:00-8:25 MCC fixing beam positions.
- 08:27 Resume 100 uA production. Lumi widths around 1400 ppm.
- 08:30-09:30 Compton rates flatlined. Laser off. Mark Dalton
reboots EPICS and has to move the GSO crystal back into its normal position.
- 08:40 Current reduced to 96 uA.
- 10:32 MCC starts injector work. Beam off.
- 10:57 Controlled access for Dave Meekins to investigate
- During controlled access Jack noticed that the left Q1 displayed
an "Invalid command" error.
Both right and left Q1's were ramped down and power cycled.
- 13:10 Controlled access completed and going back to beam permit.
- 13:30 MCC sending 100 uA into Hall A, but Compton HV tripping off.
- 14:53 Injector network switch dies and takes several injector
Injector RF phases got hosed and injector work starts all over again.
- Day shift ends with no beam and injector work in progress.
Coulombs So Far
- We are currently on Slug # 24 (IHWP IN)
- Coulombs taken in all so far: 132.0 C as of Tuesday 13:00.
- Please always check online hists and compare them to the
- If you find something strange or wrong, please HALOG it so
that the expert can check it.
- Target Operator, please always check target PID system and
strip plots to make sure all the systems work.
- The cryo flow rate around 25.2g/s and JT value is around
55.5 since this is better for Lumi widths.
- The flow rate and JT value setting are sensitive to the outside
temperature. Hence the
settings may change from morning to evening.
- Please be patient when you adjust the JT value since it
may take a few minutes to make the cryo flow rate stable.